OpenStudy (anonymous):

The ratio of two numbers is 7 to 5. The sum of the numbers is 24. What are the numbers?

OpenStudy (anonymous):

x+y=24, x/y=7/5 Solve system of equations i guess. BUt not sure.

OpenStudy (anonymous):

okay thanks timo

OpenStudy (binary3i):

let the no be X and Y so X/Y = 7/5 X+Y=24 so X=7Y/5 substitute it in the equation Y+ 7Y/5 =24 Y=10 and get the X X=14

OpenStudy (anonymous):

thank you soo much @binary3i

OpenStudy (anonymous):

x = 14, y = 10

OpenStudy (anonymous):

yes!!! thank you guys for helping me out i appreciate that :)

OpenStudy (raden):

i was use stupid way :) 7/5 = 14/10 because 14+10=24, so the first number is 14 and the other be 10

OpenStudy (anonymous):

its okay @RadEn you mtried thanks for thank fan me guys!! <3

OpenStudy (anonymous):

*tried

OpenStudy (binary3i):

@RadEn its not stupid way. its trial and error method

OpenStudy (raden):

lol, just kidding man :)